# Split Gibbs Discrete Diffusion Posterior Sampling

### v1

@@ Abstract
We study the problem of posterior sampling in discrete-state spaces using discrete diffusion models. While posterior sampling methods for continuous diffusion models have achieved remarkable progress, analogous methods for discrete diffusion models remain challenging. In this work, we introduce a principled plug-and-play discrete diffusion posterior sampling algorithm based on split Gibbs sampling, which we call SG-DPS. Our algorithm enables reward-guided generation and solving inverse problems in discrete-state spaces. We demonstrate that SG-DPS converges to the true posterior distribution on synthetic benchmarks, and enjoys state-of-the-art posterior sampling performance on a range of benchmarks for discrete data, achieving up to 2x improved performance compared to existing baselines.
@@ 1 Introduction
We study the problem of posterior sampling with discrete diffusion models. Given a pretrained discrete diffusion model (Sohl-Dickstein et al., 2015; Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024) that models the prior distribution $p({\mathbf{x}})$, our goal is to sample from the posterior $p({\mathbf{x}}|{\mathbf{y}})$ in a plug-and-play fashion, where ${\mathbf{y}}$ represents a guiding signal of interest. This setting encompasses both solving inverse problems (e.g., if ${\mathbf{y}}$ is an incomplete measurement of ${\mathbf{x}}$), and conditional or guided generation (e.g., sampling from $p({\mathbf{x}})\exp(r({\mathbf{x}})/\alpha))/Z$ according to a reward function $r(\cdot)$).
@@
Existing works on diffusion posterior sampling (Chung et al., 2023; Song et al., 2023a; Mardani et al., 2023; Wu et al., 2024) primarily focus on continuous diffusion models operating in the Euclidean space for ${\mathbf{x}}$. These methods have achieved remarkable success in applications including natural image restorations (Kawar et al., 2022; Chung et al., 2023; Zhu et al., 2023), medical imaging (Jalal et al., 2021; Song et al., 2022; Chung & Ye, 2022) and various scientific inverse problems (Feng et al., 2023; Wu et al., 2024; Zheng et al., 2024a). However, posterior sampling with discrete diffusion models in discrete-state spaces remains a challenging problem, mainly due to not having well-defined gradients of the likelihood function $p({\mathbf{y}}|{\mathbf{x}})$. Initial work to address this gap for discrete diffusion models relies on approximating value functions (similar to reinforcement learning) for derivative-free sampling (Li et al., 2024), which can be difficult to tune for good performance, particularly when the guidance ${\mathbf{y}}$ is complicated.
@@
In this paper, we develop a rigorous plug-and-play method based on the principle of split Gibbs sampling (Vono et al., 2019), which we refer to as SG-DPS. Our method addresses the posterior sampling problem ${\mathbf{x}}\sim p({\mathbf{x}}|{\mathbf{y}})$ by introducing an auxiliary variable ${\mathbf{z}}$ along with a regularization potential function $D({\mathbf{x}},{\mathbf{z}};\rho)$ to relax the original problem. The relaxed formulation enables efficient sampling via Gibbs sampling, which alternates between a likelihood sampling step and a prior sampling step using a pretrained discrete diffusion model. As regularization increases, both variables ${\mathbf{x}}$ and ${\mathbf{z}}$ converge to the target posterior distribution.
@@
We validate SG-DPS on diverse inverse problems and reward guided generation problems that involve discrete data. We demonstrate that SG-DPS converges to the true posterior distribution on synthetic data, that SG-DPS solves discrete inverse problems accurately and efficiently on discretized image data and monophonic music data, and that SG-DPS excels at reward-guided sampling on guided DNA generation. Our method demonstrates strong sampling quality in all experiments. For instance, we achieve an $8.36\ \mathrm{dB}$ improvement of PSNR values on an MNIST XOR task, over 2×2$\times$2 × smaller Hellinger distance in music infilling, and $36\%$ higher average reward in guided DNA generation compared to existing methods.
@@ 2.1 Discrete Diffusion Models
Diffusion models (Song & Ermon, 2019, 2020; Ho et al., 2020; Song et al., 2021; Karras et al., 2022, 2024) generate data ${\mathbf{x}}\in\mathcal{X}$ by reversing a predefined forward process. When $\mathcal{X}\in\mathbb{R}^{n}$ is a Euclidean space, the forward process is typically defined to be a stochastic differential equation (Karras et al., 2022):
$\mathrm{d}{\mathbf{x}}_{t}=\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t},$ (1)
where ${\mathbf{w}}_{t}$ is a standard Wiener process and $\{\sigma_{t}\}_{t=0}^{T}$ is a predefined noise schedule. This diffusion process is reversible once the score function $s({\mathbf{x}}_{t},t):=\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is learned. Starting from ${\mathbf{x}}_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}{\bm{I}})$, we generate data following the reverse SDE:
@@
Recent works on discrete diffusion models (Sohl-Dickstein et al., 2015; Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024) extend score-based generative methods from modeling continuous distributions in Euclidean spaces to categorical distributions in discrete-state spaces. Specifically, when the data distribution lies in a finite support $\mathcal{X}=\{1,\dots,N\}$, one can evolve a family of categorical distributions $p_{t}$ over $\mathcal{X}$ following a continuous-time Markov chain over the discrete elements,
$\frac{\mathrm{d}p_{t}}{\mathrm{d}t}={\bm{Q}}_{t}p_{t},$ (3)
where $p_{0}=p_{\text{data}}$ and ${\bm{Q}}_{t}\in\mathbb{R}^{N\times N}$ are diffusion matrices with a simple stationary distribution. To reverse this continuous-time Markov chain, it suffices to learn the concrete score $s({\mathbf{x}};t):=[\frac{p_{t}(\tilde{\mathbf{x}})}{p_{t}({\mathbf{x}})}]_{\tilde{\mathbf{x}}\neq{\mathbf{x}}}$, as the reverse process is given by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}=\bar{\bm{Q}}_{T-t}p_{T-t},$ (4)
with $\bar{\bm{Q}}_{t,\tilde{\mathbf{x}}{\mathbf{x}}}=\frac{p_{t}(\tilde{\mathbf{x}})}{p_{t}({\mathbf{x}})}{\bm{Q}}_{t,{\mathbf{x}}\tilde{\mathbf{x}}}$ and $\bar{\bm{Q}}_{t,{\mathbf{x}}{\mathbf{x}}}=-\sum_{\tilde{\mathbf{x}}\neq{\mathbf{x}}}\bar{\bm{Q}}_{t,\tilde{\mathbf{x}}{\mathbf{x}}}$.
For sequential data ${\mathbf{x}}\in\mathcal{X}^{D}$, there is $N^{D}$ number of states in total. Instead of constructing an exponentially large diffusion matrix, we choose a sparse matrix that only perturbs tokens independently, i.e.,
${\bm{Q}}_{t}({\mathbf{x}},\tilde{\mathbf{x}})=\sum_{d=1}^{D}{\bm{Q}}_{t}^{d}({\mathbf{x}}_{d},\tilde{\mathbf{x}}_{d})\delta_{{\mathbf{x}}_{1:D\backslash d},\tilde{\mathbf{x}}_{1:D\backslash d}},$ (5)
so that ${\bm{Q}}_{t}$ is nonzero only when the hamming distance between ${\mathbf{x}}$ and $\tilde{\mathbf{x}}$ is no more than 1.
@@
${\bm{Q}}_{t}=\dot{\sigma}_{t}{\bm{Q}}^{\mathrm{uniform}}=\dot{\sigma}_{t}\left(\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-{\bm{I}}\right),$ (6)
@@
$p_{t}$ $=\exp(\int_{0}^{t}{\bm{Q}}_{\tau}\mathrm{d}\tau)p_{0}=\exp(\sigma_{t}{\bm{Q}}^{\mathrm{uniform}})p_{0}$
$=\Big{(}e^{-\sigma_{t}}{\bm{I}}+(1-e^{-\sigma_{t}})\frac{\mathbf{1}\mathbf{1}^{T}}{N}\Big{)}p_{0}.$ (7)
When ${\mathbf{x}}\in\mathcal{X}^{D}$ is a sequence of length $D$,
$p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\ \propto\ \beta_{t}^{d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}(1-\beta_{t})^{D-d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}.$ (8)
$d(\cdot,\cdot)$ is the Hamming distance between two sequences, and $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$.
@@ 2.2 Diffusion Posterior Sampling Methods
Diffusion posterior sampling is a class of methods that draw samples from $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$ using a diffusion model trained on the prior $p({\mathbf{x}})$. This technique has been applied in two settings with slightly different formulations: solving inverse problems and loss/reward guided generation.
@@
On the other hand, reward-guided generation (Song et al., 2023b; Huang et al., 2024; Nisonoff et al., 2024) focuses on generating samples that achieve higher rewards according to a given reward function $r(\cdot)$. Here, the likelihood is modeled as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(r({\mathbf{x}})/\alpha)$, where $\alpha>0$ is a temperature parameter that controls the strength of reward weighting.
@@
Existing diffusion posterior sampling methods leverage (continuous) diffusion models as plug-and-play priors and modifies Equation 2 to generate samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$. These methods can be broadly classified into four categories (Daras et al., 2024; Zheng et al., 2024b).
Guidance-based methods (Chung et al., 2023; Song et al., 2023b, a; Kawar et al., 2022; Wang et al., 2023; Zheng et al., 2024a) perform posterior sampling by adding an additional likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})$ to the score function in Equation 2, and modifies the reverse SDE as follows:
$\mathrm{d}{\mathbf{x}}_{t}$ $=-2\dot{\sigma}_{t}\sigma_{t}(\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})+\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t}))\mathrm{d}t$
$\hskip 130.0002pt+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (9)
@@
The intractable likelihood score term is approximated with various assumptions. However, these methods cannot be directly applied to discrete posterior sampling problems, where $\log p({\mathbf{y}}|{\mathbf{x}})$ is defined only on a finite support.
@@
Sequential Monte Carlo (SMC) methods sample batches of particles iteratively from a series of distributions, which converge to the posterior distribution in limit. This class of methods (Cardoso et al., 2024; Dou & Song, 2024) exploits the sequential diffusion sampling process to sample from the posterior distribution. Similar ideas have been applied in Li et al. (2024) for reward-guided generation using discrete diffusion models, which rely on approximating a soft value function at each iteration. We compare our method to this approach in our experiments.
@@
Variational methods (Mardani et al., 2023; Feng et al., 2023) solve inverse problems by approximating the posterior distribution with a parameterized distribution $q_{\bm{\theta}}$ and optimizing its Kullback-Leibler (KL) divergence with respect to the posterior. However, how to extend these methods to discrete diffusion models remains largely unexplored.
@@
Variable splitting methods (Wu et al., 2024; Zhang et al., 2024; Zhu et al., 2023; Song et al., 2024; Xu & Chi, 2024) decompose posterior sampling into two simpler alternating steps. A representative of this category is the split Gibbs sampler (see Section 3), where the variable ${\mathbf{x}}_{t}$ is split into two random variables that both eventually converge to the posterior distribution. These methods are primarily designed for continuous diffusion models with Gaussian kernels, leveraging Langevin dynamics or optimization in Euclidean space as subroutines. Murata et al. (2024) extends this category to VQ-diffusion models in the discrete latent space using the Gumbel-softmax reparameterization trick. However, this approach is limited to discrete latent spaces with an underlying continuous embedding due to its reliance on softmax dequantization. In this work, we generalize the split Gibbs sampler to posterior sampling in categorical data by further exploiting the unique properties of discrete diffusion models.
@@
Figure 1: An illustration of our method on inpainting discretized MNIST. Our method implements the Split Gibbs sampler, which alternates between likelihood sampling steps and prior sampling steps. Likelihood sampling steps enforce data consistency with regard to measurement ${\mathbf{y}}$, while prior sampling steps involve denoising uniform noise from $z^{(k)}$ by a pretrained discrete diffusion model starting from noise level $\rho_{k}$. Two variables converge to a sample from the posterior distribution as the noise level $\rho_{k}$ reduces to zero.

### v2

@@ Abstract
We study the problem of posterior sampling in discrete-state spaces using discrete diffusion models. While posterior sampling methods for continuous diffusion models have achieved remarkable progress, analogous methods for discrete diffusion models remain challenging. In this work, we introduce a principled plug-and-play discrete diffusion posterior sampling algorithm based on split Gibbs sampling, which we call SGDD. Our algorithm enables reward-guided generation and solving inverse problems in discrete-state spaces. We demonstrate the convergence of SGDD to the target posterior distribution and verify this through controlled experiments on synthetic benchmarks. Our method enjoys state-of-the-art posterior sampling performance on a range of benchmarks for discrete data, including DNA sequence design, discrete image inverse problems, and music infilling, achieving more than $30\%$ improved performance compared to existing baselines. Our code is available at SGDD.
@@ 1 Introduction
We study the problem of posterior sampling in finite spaces with discrete diffusion models. Given a pretrained discrete diffusion model [33, 1, 4, 29] that models the prior distribution $p({\mathbf{x}})$, our goal is to sample from the posterior $p({\mathbf{x}}|{\mathbf{y}})$ in a plug-and-play fashion, where ${\mathbf{y}}$ represents a guiding signal of interest. This setting encompasses both solving inverse problems (e.g., if ${\mathbf{y}}$ is an incomplete measurement of ${\mathbf{x}}$), and conditional or guided generation (e.g., sampling from $p({\mathbf{x}})\exp(\beta r({\mathbf{x}})))/Z$ according to a reward function $r(\cdot)$).
@@
Existing works on diffusion posterior sampling [7, 35, 30, 48] primarily focus on continuous diffusion models operating in the Euclidean space for ${\mathbf{x}}$. These methods have achieved remarkable success in applications including natural image restorations [24, 7, 54], medical imaging [20, 40, 6], and various scientific inverse problems [14, 48, 52]. However, posterior sampling with discrete diffusion models in discrete-state spaces remains a challenging problem, mainly due to not having well-defined gradients of the likelihood function $p({\mathbf{y}}|{\mathbf{x}})$. Initial work to address this gap for discrete diffusion models relies on approximating value functions (similar to reinforcement learning) for derivative-free sampling [27], which can be difficult to tune for good performance, particularly when the guidance ${\mathbf{y}}$ is complicated.
@@
In this paper, we develop a rigorous plug-and-play method based on the principle of split Gibbs sampling [44], which we refer to as SGDD. Our method addresses the posterior sampling problem ${\mathbf{x}}\sim p({\mathbf{x}}|{\mathbf{y}})$ by introducing an auxiliary variable ${\mathbf{z}}$ along with a regularization potential function $D({\mathbf{x}},{\mathbf{z}};\eta)$ to relax the original problem. The relaxed formulation enables efficient sampling via Gibbs sampling, which alternates between a likelihood sampling step and a prior sampling step using a pretrained discrete diffusion model. As regularization increases, both variables ${\mathbf{x}}$ and ${\mathbf{z}}$ converge to the target posterior distribution. Our method generalizes split Gibbs samplers with $\ell_{2}$ regularizers to any potential functions with certain limiting conditions, making it compatible with discrete diffusion models by carefully designing the potential function. We provide a rigorous stationary guarantee on the convergence of SGDD, with error bounds that account for imperfect score function and discretization errors.
@@
We validate SGDD on diverse inverse problems and reward-guided generation problems that involve discrete data. We demonstrate that SGDD converges to the true posterior distribution on synthetic data, that SGDD solves discrete inverse problems accurately and efficiently on discretized image data and monophonic music data, and that SGDD excels at reward-guided sampling on guided DNA generation. Our method demonstrates strong sampling quality in all experiments. For instance, we achieve a $42\%$ higher median activity in enhancer DNA design, an $8.36\ \mathrm{dB}$ improvement of PSNR values in solving the XOR inverse problem on MNIST, and over 2×2$\times$2 × smaller Hellinger distance in music infilling, compared to existing methods.
@@ 2.1 Discrete Diffusion Models
Diffusion models [37, 38, 18, 39, 22, 23] generate data ${\mathbf{x}}\in\mathcal{X}$ by reversing a predefined forward process. When $\mathcal{X}\in\mathbb{R}^{n}$ is a Euclidean space, the forward process is typically defined to be a stochastic differential equation [22]: $\mathrm{d}{\mathbf{x}}_{t}=\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t},$ where ${\mathbf{w}}_{t}$ is a standard Wiener process and $\{\sigma_{t}\}_{t=0}^{T}$ is a predefined noise schedule. This diffusion process is reversible once the score function $s({\mathbf{x}}_{t},t):=\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is learned. Starting from ${\mathbf{x}}_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}{\bm{I}})$, we generate data following the reverse SDE:
@@
Recent works on discrete diffusion models [33, 1, 4, 29] extend score-based generative methods from modeling continuous distributions in Euclidean spaces to categorical distributions in discrete-state spaces. Specifically, when the data distribution lies in a finite support $\mathcal{X}=\{1,\dots,N\}$, one can evolve a family of categorical distributions $p_{t}$ over $\mathcal{X}$ following a continuous-time Markov chain over the discrete elements,
$\partial_{t}p_{t}={\bm{Q}}_{t}^{\mathsf{fw}}p_{t},$ (2)
where $p_{0}=p_{\text{data}}$ and ${\bm{Q}}^{\mathsf{fw}}_{t}\in\mathbb{R}^{N\times N}$ are diffusion matrices with a simple stationary distribution. To reverse this continuous-time Markov chain, it suffices to learn the concrete score $s({\mathbf{x}};t):=[\frac{p_{t}(\tilde{\mathbf{x}})}{p_{t}({\mathbf{x}})}]_{\tilde{\mathbf{x}}\neq{\mathbf{x}}}$, as the reverse process is given by
$\partial_{t}p_{T-t}={\bm{Q}}_{t}p_{T-t}$ (3)
with ${\bm{Q}}_{t}^{[i,j]}=\frac{p_{T-t}({\mathbf{x}}_{i})}{p_{T-t}({\mathbf{x}}_{j})}{\bm{Q}}_{T-t}^{\mathsf{fw}[j,i]}$ and ${\bm{Q}}_{t}^{[i,i]}=-\sum_{{\mathbf{x}}_{j}\neq{\mathbf{x}}_{i}}{\bm{Q}}_{t}^{[j,i]}$.
For sequential data ${\mathbf{x}}\in\mathcal{X}^{D}$, there are $N^{D}$ states in total. Instead of constructing an exponentially large diffusion matrix, we use a sparse matrix ${\bm{Q}}^{\mathsf{fw}}_{t}$ that perturbs tokens independently in each dimension [29].
@@
${\bm{Q}}_{t}^{\mathsf{fw}}=\dot{\sigma}_{t}{\bm{Q}}^{\mathrm{uniform}}=\dot{\sigma}_{t}(\mathbf{1}\mathbf{1}^{T}/N-{\bm{I}}),$ (4)
@@
$p_{t}$ $=\exp\left(\int_{0}^{t}{\bm{Q}}_{\tau}^{\mathsf{fw}}\mathrm{d}\tau\right)p_{0}=\exp(\sigma_{t}{\bm{Q}}^{\mathrm{uniform}})p_{0}=\Big{(}e^{-\sigma_{t}}{\bm{I}}+(1-e^{-\sigma_{t}})\frac{\mathbf{1}\mathbf{1}^{T}}{N}\Big{)}p_{0}.$ (5)
When ${\mathbf{x}}\in\mathcal{X}^{D}$ is a sequence of length $D$, we have $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\ \propto\ \beta_{t}^{d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}(1-\beta_{t})^{D-d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}$, where $d(\cdot,\cdot)$ is the Hamming distance between two sequences, and $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$.
@@ 2.2 Diffusion Posterior Sampling Methods
Diffusion posterior sampling is a class of methods that draw samples from $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$ using a diffusion model trained on the prior $p({\mathbf{x}})$. This technique has been applied in two settings with slightly different formulations: solving inverse problems and reward guided generation.
@@
On the other hand, reward-guided generation [36, 19, 32] focuses on generating samples that achieve higher rewards according to a given reward function $r(\cdot)$. Here, the likelihood is modeled as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(\beta r({\mathbf{x}}))$, where $\beta>0$ is a parameter that controls the strength of reward weighting.
@@
Existing diffusion posterior sampling methods leverage (continuous) diffusion models as plug-and-play priors and modify Eq. 1 to generate samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$. These methods can be broadly classified into four categories [9, 53].
@@
Guidance-based methods [7, 36, 35, 24, 46, 52] perform posterior sampling by adding an additional likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})$ to the score function in Eq. 1, and modifies the reverse SDE as: $\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}(\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})+\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t}))\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ The intractable likelihood score term is approximated with various assumptions. However, these methods cannot be directly applied to discrete posterior sampling problems, where $\log p({\mathbf{y}}|{\mathbf{x}})$ is defined only on a finite support and thus does not have such gradient information.
@@
Sequential Monte Carlo (SMC) methods sample batches of particles iteratively from a series of distributions, which converge to the posterior distribution in the limit. This class of methods [5, 12, 26, 47] exploits the sequential diffusion sampling process to sample from the posterior distribution. Similar ideas have been applied in Li et al. [27] for reward-guided generation using discrete diffusion models, which rely on approximating a soft value function at each iteration.
@@
Variational methods [30, 14] solve inverse problems by approximating the posterior distribution with a parameterized distribution $q_{\bm{\theta}}$ and optimizing its Kullback-Leibler (KL) divergence with respect to the posterior. However, how to extend these methods to discrete diffusion models remains largely unexplored.
@@
Variable splitting methods [48, 51, 54, 34, 49] decompose posterior sampling into two simpler alternating steps. A representative of this category is the split Gibbs sampler (see Section 3), where the variable ${\mathbf{x}}_{t}$ is split into two random variables that both eventually converge to the posterior distribution. These methods are primarily designed for continuous diffusion models with Gaussian kernels, leveraging Langevin dynamics or optimization in Euclidean space as subroutines. Murata et al. [31] extends this category to VQ-diffusion models in the discrete latent space using the Gumbel-softmax reparameterization trick. However, this approach is limited to discrete latent spaces with an underlying continuous embedding due to its reliance on softmax dequantization. In this work, we generalize the split Gibbs sampler to posterior sampling in categorical data by further exploiting the unique properties of discrete diffusion models.
@@
Figure 1: An illustration of our method on inpainting discretized MNIST. Our method implements the Split Gibbs sampler, which alternates between likelihood sampling steps and prior sampling steps. Likelihood sampling steps enforce data consistency with regard to measurement ${\mathbf{y}}$, while prior sampling steps involve denoising uniform noise from $z^{(k)}$ by a pretrained discrete diffusion model starting from noise level $\eta_{k}$. Two variables converge to a sample from the posterior distribution as the noise level $\eta_{k}$ reduces to zero.